# Overlap reduction functions for pulsar timing arrays and astrometry

### v1

@@ III.2.2 Spin-1 GWs
$\bar{E}(\hat{n}|\hat{k})$ $=\tilde{h}_{x}\frac{(\sin 2\theta+\sin\theta)\cos\phi}{1+\cos\theta}+\tilde{h}_{y}\frac{(\sin 2\theta+\sin\theta)\sin\phi}{1+\cos\theta}$
@@
$\bar{B}(\hat{n}|\hat{k})$ $=\tilde{h}_{x}\frac{\sin 2\theta\sin\phi}{2(1+\cos\theta)}-\tilde{h}_{y}\frac{\sin 2\theta\cos\phi}{2(1+\cos\theta)}$
@@
$\tilde{E}_{f,\ell m}(\hat{k})$ $=\frac{1}{\sqrt{\ell(\ell+1)}}\int^{2\pi}_{0}\mathrm{d}\phi\int^{\theta}_{0}\mathrm{d}\theta\sin\theta\,Y^{*}_{\ell m}(\theta,\phi)\bar{E}(\hat{n}|\hat{k})$
@@
$\tilde{B}_{f,\ell m}(\hat{k})$ $=\frac{1}{\sqrt{\ell(\ell+1)}}\int^{2\pi}_{0}\mathrm{d}\phi\int^{\theta}_{0}\mathrm{d}\theta\sin\theta\,Y^{*}_{\ell m}(\theta,\phi)\bar{B}(\hat{n}|\hat{k})$
@@
$=2E^{v}_{\ell}B^{v}_{\ell^{\prime}}I^{v}_{0}\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[c^{v,I}_{LM}(1-X^{L}_{\ell\ell^{\prime}})-c^{v,V}_{LM}X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -1&1&0\\ \end{pmatrix}+\left[-c^{v,E}_{LM}(1-X^{L}_{\ell\ell^{\prime}})-ic^{v,B}_{LM}X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 1&1&-2\\ \end{pmatrix}\right\},$ (74)
@@
$=2B^{v*}_{\ell}E^{v}_{\ell^{\prime}}I^{v}_{0}\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[c^{v,I}_{LM}(1-X^{L}_{\ell\ell^{\prime}})-c^{v,V}_{LM}X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -1&1&0\\ \end{pmatrix}+\left[c^{v,E}_{LM}(1-X^{L}_{\ell\ell^{\prime}})+ic^{v,B}_{LM}X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 1&1&-2\\ \end{pmatrix}\right\},$ (75)
@@
$=2B^{v*}_{\ell}B^{v}_{\ell^{\prime}}I^{v}_{0}\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[-c^{v,I}_{LM}X^{L}_{\ell\ell^{\prime}}+c^{v,V}_{LM}(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -1&1&0\\ \end{pmatrix}+\left[-c^{v,E}_{LM}X^{L}_{\ell\ell^{\prime}}-ic^{v,B}_{LM}(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 1&1&-2\\ \end{pmatrix}\right\},$ (76)
@@ III.3 Cross-correlation between redshift response and deflection fields
$=\sum_{p=t,v}\sum_{X}\sum_{S}i^{X}\left[I^{p}_{0}\sum_{L,M}c^{p,X}_{LM}\Gamma^{p,X,zS}_{LM,i}(\hat{n}_{a},\hat{n}_{b})\right],$ (78)
@@
$=\sum_{p=t,v}\sum_{X}\sum_{S}i^{X}\left[I^{p}_{0}\sum_{L,M}c^{p,X}_{LM}\Gamma^{p,X,Sz}_{LM,i}(\hat{n}_{a},\hat{n}_{b})\right],$ (79)
@@
Figures 5 and 6 show the $\theta$ dependence of $\Gamma^{t,I,Ez}_{LM}$ and $\Gamma^{v,I,Ez}_{LM}$, respectively.
@@ IV Conclusion
We have checked that our results for PTA timing residuals for an anisotropic, unpolarized, spin-2 and spin-1 GW background agree with Ref. [50] up to $L\leq 3$. (While we have verified that our results are consistent for $L\leq 3$, we have only plotted up to $L\leq 1$.) For an anisotropic, polarized (either circular or linear) GW background, our results agree with those in Ref. [45]. For astrometric angular deflections with $L=M=0$ with spin-2 and spin-1 GW background, our results correspond with those in Refs. [24, 26, 46, 47, 25]. The results presented here for astrometry with $L\geq 1$ are new. Note that we show only a limited number of plots in this paper and refer to our Github page for the other ORFs.
@@ Acknowledgements
$f(\hat{n}|\hat{k})=\sum_{\ell m}\tilde{f}_{\ell m}(\hat{k})Y_{\ell m}(\hat{n}|\hat{k}),$ (94)
where $\hat{n}$ is defined in the coordinates with $\hat{k}$ aligned along the $z$-axis. We now move to other coordinates where $\hat{\Omega}(\neq\hat{k})$ is aligned along the $z$-axis. Even if the physical observation direction is fixed, the coordinate transformation changes the observation direction vector as $\hat{n}\rightarrow\hat{n}^{\prime}$. That is, we have
$f(\hat{n}|\hat{k})=f(\hat{n}^{\prime}|\hat{\Omega})=\sum_{\ell m}\tilde{f}_{\ell m}(\hat{k})Y_{\ell m}(\hat{n}|\hat{k})=\sum_{\ell m}\tilde{f}_{\ell m}(\hat{\Omega})Y_{\ell m}(\hat{n}^{\prime}|\hat{\Omega}).$ (95)
We here relate $Y_{\ell m}(\hat{n}|\hat{k})$ and $Y_{\ell m}(\hat{n}^{\prime}|\hat{\Omega})$ using Wigner $D$-matrix. We here define $R$ as
$\hat{k}^{i}=R^{ij}\hat{\Omega}_{j}.$ (96)
The Euler angles are $\alpha=\phi_{k}$ and $\beta=\theta_{k}$, where $\hat{k}=(\sin\theta_{k}\cos\phi_{k},\sin\theta_{k}\sin\phi_{k},\cos\theta_{k})$ in the $\hat{\Omega}\parallel\hat{z}$ coordinates. The final Euler angle $\gamma$ is arbitrary and we take $\gamma=0$ in this work, which corresponds to make the polarization of ‘$+$’, ‘$\times$’, ‘$x$’, and ‘$y$’ aligned with $\hat{\theta}$ and $\hat{\phi}$ directions [40]. Then, we obtain
$\hat{n}^{i}=(R^{-1})^{ij}\hat{n}^{\prime}_{j}.$ (97)
@@
$Y_{\ell m}(\hat{n}|\hat{k})$ $=\sum_{m^{\prime}}D^{(\ell)*}_{mm^{\prime}}(R^{-1})Y_{\ell m^{\prime}}(\hat{n}^{\prime}|\hat{\Omega})$
$=\sum_{m^{\prime}}Y_{\ell m^{\prime}}(\hat{n}^{\prime}|\hat{\Omega})D^{(\ell)}_{m^{\prime}m}(R).$ (98)
@@
$f(\hat{n}^{\prime}|\hat{\Omega})=\sum_{\ell}\sum_{mm^{\prime}}Y_{\ell m}(\hat{n}^{\prime}|\hat{\Omega})D^{(\ell)}_{mm^{\prime}}(\phi_{k},\theta_{k},0)\tilde{f}_{\ell m^{\prime}}(\hat{k}).$ (99)

### v2

@@ III.2.2 Spin-1 GWs
$\bar{E}(\hat{n},\hat{k})$ $=\tilde{h}_{x}\frac{(\sin 2\theta+\sin\theta)\cos\phi}{1+\cos\theta}+\tilde{h}_{y}\frac{(\sin 2\theta+\sin\theta)\sin\phi}{1+\cos\theta}$
@@
$\bar{B}(\hat{n},\hat{k})$ $=\tilde{h}_{x}\frac{\sin 2\theta\sin\phi}{2(1+\cos\theta)}-\tilde{h}_{y}\frac{\sin 2\theta\cos\phi}{2(1+\cos\theta)}$
@@
$\tilde{E}_{f,\ell m}(\hat{k})$ $=\frac{1}{\sqrt{\ell(\ell+1)}}\int^{2\pi}_{0}\mathrm{d}\phi\int^{\theta}_{0}\mathrm{d}\theta\sin\theta\,Y^{*}_{\ell m}(\theta,\phi)\bar{E}(\hat{n},\hat{k})$
@@
$\tilde{B}_{f,\ell m}(\hat{k})$ $=\frac{1}{\sqrt{\ell(\ell+1)}}\int^{2\pi}_{0}\mathrm{d}\phi\int^{\theta}_{0}\mathrm{d}\theta\sin\theta\,Y^{*}_{\ell m}(\theta,\phi)\bar{B}(\hat{n},\hat{k})$
@@
$=2E^{v}_{\ell}B^{v}_{\ell^{\prime}}I^{v}_{0}(f)\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[c^{v,I}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})-c^{v,V}_{LM}(f)X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -1&1&0\\ \end{pmatrix}+\left[-c^{v,E}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})-ic^{v,B}_{LM}(f)X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 1&1&-2\\ \end{pmatrix}\right\},$ (74)
@@
$=2B^{v*}_{\ell}E^{v}_{\ell^{\prime}}I^{v}_{0}(f)\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[c^{v,I}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})-c^{v,V}_{LM}(f)X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -1&1&0\\ \end{pmatrix}+\left[c^{v,E}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})+ic^{v,B}_{LM}(f)X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 1&1&-2\\ \end{pmatrix}\right\},$ (75)
@@
$=2B^{v*}_{\ell}B^{v}_{\ell^{\prime}}I^{v}_{0}(f)\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[-c^{v,I}_{LM}(f)X^{L}_{\ell\ell^{\prime}}+c^{v,V}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -1&1&0\\ \end{pmatrix}+\left[-c^{v,E}_{LM}(f)X^{L}_{\ell\ell^{\prime}}-ic^{v,B}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 1&1&-2\\ \end{pmatrix}\right\},$ (76)
@@ III.3 Cross-correlation between redshift response and deflection fields
$=\sum_{p=t,v}\sum_{X}\sum_{S}i^{X}\left[I^{p}_{0}(f)\sum_{L,M}c^{p,X}_{LM}(f)\Gamma^{p,X,zS}_{LM,i}(\hat{n}_{a},\hat{n}_{b})\right],$ (78)
@@
$=\sum_{p=t,v}\sum_{X}\sum_{S}i^{X}\left[I^{p}_{0}(f)\sum_{L,M}c^{p,X}_{LM}(f)\Gamma^{p,X,Sz}_{LM,i}(\hat{n}_{a},\hat{n}_{b})\right],$ (79)
@@
Figures 7 and 8 show the $\theta$ dependence of $\Gamma^{t,I,Ez}_{LM}$ and $\Gamma^{v,I,Ez}_{LM}$, respectively.
@@ IV Conclusion
We have checked that our results for PTA timing residuals for an anisotropic, unpolarized, spin-2 and spin-1 GW background agree with Ref. [50] up to $L\leq 3$. (While we have verified that our results are consistent for $L\leq 3$, we have only plotted up to $L\leq 1$.) For an anisotropic, polarized (either circular or linear) GW background, our results agree with those in Ref. [45]. For astrometric angular deflections with $L=M=0$ with spin-2 and spin-1 GW background, our results correspond with those in Refs. [24, 26, 46, 47, 25]. The results presented here for astrometry with $L\geq 1$ are new. Note that we show only a limited number of plots in this paper and refer to our Github page (Jupyter Notebook with Python code) for the other ORFs.
@@ Acknowledgements
$f_{\hat{k}}(\hat{n})=\sum_{\ell m}\tilde{f}_{\ell m}(\hat{k})Y_{\ell m}(\hat{n}),$ (94)
where the subscript $\hat{k}$ means that $\hat{n}$ is defined in the coordinates with $\hat{k}$ aligned along the $z$-axis. We now move to other coordinates where $\hat{\Omega}(\neq\hat{k})$ is aligned along the $z$-axis. Even if the physical observation direction is fixed, the coordinate transformation changes the observation direction vector as $\hat{n}\rightarrow\hat{n}^{\prime}$. That is, we have
$f_{\hat{k}}(\hat{n})=f_{\hat{\Omega}}(\hat{n}^{\prime})=\sum_{\ell m}\tilde{f}_{\ell m}(\hat{k})Y_{\ell m}(\hat{n})=\sum_{\ell m}\tilde{f}_{\ell m}(\hat{\Omega})Y_{\ell m}(\hat{n}^{\prime}).$ (95)
We here relate $Y_{\ell m}(\hat{n})$ and $Y_{\ell m}(\hat{n}^{\prime})$ using Wigner $D$-matrix. We here define $R$ as
$\hat{\Omega}^{i}=R^{ij}\hat{k}_{j}.$ (96)
The Euler angles are $\alpha=\phi_{\Omega}$ and $\beta=\theta_{\Omega}$, where $\hat{\Omega}=(\sin\theta_{\Omega}\cos\phi_{\Omega},\sin\theta_{\Omega}\sin\phi_{\Omega},\cos\theta_{\Omega})$ in the $\hat{k}\parallel\hat{z}$ coordinates. The final Euler angle $\gamma$ is arbitrary and we take $\gamma=0$ in this work, which corresponds to make the polarization of ‘$+$’, ‘$\times$’, ‘$x$’, and ‘$y$’ aligned with $\hat{\theta}$ and $\hat{\phi}$ directions [40]. Then, we obtain
${\hat{n}}^{\prime\,i}=(R^{-1})^{ij}\hat{n}_{j}.$ (97)
@@
$Y_{\ell m}(\hat{n}^{\prime})$ $=\sum_{m^{\prime}}D^{(\ell)*}_{mm^{\prime}}(R^{-1})Y_{\ell m^{\prime}}(\hat{n})$
$=\sum_{m^{\prime}}Y_{\ell m^{\prime}}(\hat{n})D^{(\ell)}_{m^{\prime}m}(R).$ (98)
@@
$f_{\hat{k}}(\hat{n})=\sum_{\ell}\sum_{mm^{\prime}}Y_{\ell m}(\hat{n})D^{(\ell)}_{mm^{\prime}}(\phi_{\Omega},\theta_{\Omega},0)\tilde{f}_{\ell m^{\prime}}(\hat{\Omega}).$ (99)